Co (s) | Co2+ (aq) || Au3+ (aq) | Au (s)

Q3\ For the previous cell, determine the concentration of the cobalt ion when a cell voltage of 1.76 V is obtained using a 0.761 M gold (III) nitrate solution.  

Q4\ Calculate the free energy for the nonstandard cell in the previous problem. Show all work. Be mindful of signs, sig figs, and units

Q5\ Consider a galvanic cell consisting of the following two redox couples: Sn2+ (aq, 0.010 M) + 2e- --> Sn (s) Eo = -0.14 V Cs+ (aq, 0.010 M) + e- --> Cs (s) Eo = -2.92 V

a. Write the equation for the half-reaction at the cathode

b. Write the equation for the half-reaction at the anode

c. Write a shorthand cell notation version of this battery

d. Calculate the standard cell potential for this battery in volts. Include the correct number of sig figs in your answer.
Answer #1

CO - > Co tot ze Au 3t +3é Au 3x2 300 + 2 Aut - 3 Co2+ + 2 Au EColat te ruste) -0.0591 log [comme Austy 1.76 = (0,28 +1.50) -04. AC = - nFEO = - 64 96500 × 1.78 = -1030620 Jmolt - AG3 AG AC = A + RI and - = -1030620 + 8.314 x298 en cou)? - SAU2 = -10as. a) snt + z ł SA 6) cs cst té c) colest 11 snt Isn d) E = € sest + Espalion - 0.059) log fesy 2 E = 2.92 + (-0.14) - 6 (0.